what is the ph of 0.50M solution of a weak base the kb value for this base is 2.5*10^-6
Answer #1

B dissociates as:

B +H2O -----> BH+ + OH-

0.5 0 0

0.5-x x x

Kb = [BH+][OH-]/[B]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((2.5*10^-6)*0.5) = 1.118*10^-3

since c is much greater than x, our assumption is correct

so, x = 1.118*10^-3 M

So, [OH-] = x = 1.118*10^-3 M

use:

pOH = -log [OH-]

= -log (1.118*10^-3)

= 2.9515

use:

PH = 14 - pOH

= 14 - 2.9515

= 11.0485

Answer: 11.05